OpenStudy (anonymous):

Math help;// Write an EQUATION for the problem and SOLVE. 1. The Drama Club at at Smith Middle School is selling cookie dough in order to raise money for costumes. If each tub or cookie dough costs $4.75, how many tubs must each member sell to make $570.00? 2. Nicholas bicycled 15.8 kilometers each day for 18 days last month. How many kilometers did Nicholas bicycle last month?

OpenStudy (swolesammy):

1. is 120 tubs

OpenStudy (swolesammy):

2. 284.4 km

OpenStudy (thread252):

1=120 2=284.4

OpenStudy (anonymous):

570 = 4.75 x N Where N := the number of tubs that must be sold If 570 = 4.75 x N Then N = 570 / 4.75 So N = 120 *The only issue with this is the question asks how many each member must sell, and we don't know how many members there are. 120 tubs must be sold in total.

OpenStudy (anonymous):

15.8 x 18 = N Where N := the total number of kilometres cycled N = 15.8 x 18 N = 284.4 km